OpenStudy (anonymous):

The line of symmetry for the quadratic equation y = ax^2 + 8x - 3 is x = 4. What is the value of "a"? A.) -2 B.) -1 C.) 1

jimthompson5910 (jim_thompson5910):

Axis of symmetry x = -b/(2a) 4 = -8/(2a) 4*2a = -8 ... ... ... a = ???

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank ya sir

jimthompson5910 (jim_thompson5910):

you're welcome